Translate to a system of equations and solve.
Drew burned 1,500 calories Friday playing one hour of basketball and canoeing for two hours. Saturday he spent two hours playing basketball and three hours canoeing and burned 2,600 calories. How many calories did he burn per hour when playing basketball? How many calories did he burn per hour when canoeing?
basketball ____ calories per hour
canoeing _____ calories per hour

Answer by Boreal(15235)

1B+2C=1500, where B is basketball and C canoeing
2B+3C=2600
subtract twice first from second to eliminate B
-2B-4C=-3000
-C=-400
C=400 cal/hr canoeing
B=700 cal/hr basketball